OpenStudy (anonymous):

Joe earned $300 on an amount deposited into his bank account for 3 years at a 2% annual simple interest rate. How much did Joe deposit?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Hmm.

TheSmartOne (thesmartone):

Do you know the simple interest formula?

OpenStudy (anonymous):

l=prt

OpenStudy (anonymous):

i=interest p=principal r=rate t =time

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

So what is t=? p is what we are finding i=? r=? They gave us i, r, and t in the question.

OpenStudy (anonymous):

i=interest p=principal r=rate t =time

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Yes but what are the numbers?

TheSmartOne (thesmartone):

lol

OpenStudy (anonymous):

t=3 years r=2% i =$300

OpenStudy (anonymous):

@t

TheSmartOne (thesmartone):

lol Now do you know that r has to be in decimal? So what is 2% in decimal?

OpenStudy (anonymous):

0.02

TheSmartOne (thesmartone):

Correct so now back to the main equation. We just plug everything back in to it and slve for P :)

OpenStudy (anonymous):

i still dont get it

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @morganlauren11 l=prt \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @morganlauren11 t=3 years r=2% i =$300 \(\color{blue}{\text{End of Quote}}\) And just changing r= .02 Solve for P now after you substitute all the numbers :)

OpenStudy (anonymous):

$300=p0.023 i have this but how do i find p

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

So we have \[300= p \times .02 \times 3\]

OpenStudy (anonymous):

.02 x 3 =.06

TheSmartOne (thesmartone):

Ok correct! Now we will get \[300=.06p \]

OpenStudy (anonymous):

wouldnt it then be 300 divided by .06

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Yes exactly! So what is \[\frac{ 300}{ .06 }=? \]

OpenStudy (anonymous):

0.0002

TheSmartOne (thesmartone):

Are you sure?

OpenStudy (anonymous):

either that or 5000

OpenStudy (anonymous):

sorry its 2am and im trying to finish school so i can go to sleep

TheSmartOne (thesmartone):

lol 5000 is correct!

OpenStudy (anonymous):

oh great thank you:)

TheSmartOne (thesmartone):

Np :)